The Minnesota Multiphasic Personality Inventory (MMPI) is one of the most frequently used personality tests in psychology. The current standardized version, the MMPI-2, is used by trained professionals to assist in identifying personality structure and psychopathology. A forensic psychologist is interested in whether she can identify potential serial rapists by their MMPI profiles. She is particularly interested in the Psychopathic Deviate and Infrequency scales of the MMPI. She interviews 12 people who have already been identified as serial rapists. Their scores on the Psychopathic Deviate scale are recorded here:

51 60 35 69 32 47 66 43 58 28 55

44

Which of the following is a possible value for the range?

41

49

20

59

Their scores on the Infrequency scale are given here:

33 60 26 37 22 87 76 45 55 41 48 19

Which of the following is a possible value for the range?

28

68

46

58
Answer #1

#1.
Range = max - min
max = 69
min = 28

Range = 69 - 28 = 41

#2.
max = 87
min = 19

Range = 87 - 19 = 68